OpenStudy (jylarai):

A sack of fruit contains 4 pears, 6 peaches and 5 mangos. Kevin randomly chooses a piece of fruit without putting it back. Then Mark randomly chooses a piece of fruit. What is the probabilty that Kevin chooses a mango and Mark chooses a peach. A.2/15 B.1/7 C .2/5 D. 1/3

OpenStudy (anonymous):

B.

OpenStudy (jylarai):

Can you explain how you got B so I can show my work

OpenStudy (anonymous):

Kevin choosing mango chance: (5/4)+6+5=1/3 Mark choosing peach chance: (6/4)+5+5=3/7 (3/7)(1/3)=1/7

OpenStudy (jylarai):

Oh thank you